In an enzyme reaction experiment, the concentration of the substrate S

S is varied and the velocity measured for each concentration, giving the following data.

[S](mM)= 0.2, 0.4, 0.6, 0.8, 1.0

V(μM/min)= 0.38, 0.62, 0.79, 0.91, 1

By sketching the Lineweaver-Burk plot of 1/V against 1/[S] determine an approximation of

Km and Vmax. No need to sketch.

Which of the following is the best approximation of Km and the maximum velocity Vmax?


1
Expert's answer
2021-10-27T05:40:01-0400

0.2 × 1= 0.2

0.4×0.91= 0.364

0.6×0.79=0.474

0.8×0.62= 0.496

0.38×1= 0.38

Hence the maximum is 0.8[S](mM), 0.62V(μM/min)